# Bitcoin Under Volatile Block Rewards: How Mempool Statistics Can Influence Bitcoin Mining

### v2

@@ Abstract.
This paper presents a reinforcement learning-based tool to develop mining strategies under a more realistic volatile model. We employ the Asynchronous Advantage Actor-Critic (A3C) algorithm, which efficiently handles dynamic environments, such as the Bitcoin mempool, to derive near-optimal mining strategies when interacting with an environment that models the complexity of the Bitcoin mempool. This tool enables the analysis of adversarial mining strategies, such as selfish mining and undercutting, both before and after difficulty adjustments, providing insights into the effects of mining attacks in both the short and long term.
@@ 1. Introduction
We proceed by introducing a simplified model to analyze mining strategies, in which the volatility of block rewards arises from the varying generation times of blocks. We argue that block reward volatility is not solely attributed to the infrequent arrival of whale transactions. The inherent competition among transactions to be included in the chain quickly causes blocks with longer generation times to include higher transaction fees, resulting in block reward volatility even in the absence of whale transactions. To this end, we build our simplified model upon the observed positive correlation between block generation time and its fee. Compared to the WeRLman environment, our simplified environment includes a greater number of blocks with varying fee levels, allowing us to analyze more fine-grained strategies. We employ a Markov Decision Process (MDP)-based tool to analyze the profitability of mining strategies under our simplified volatile model. Due to the limited number of states in the simplified model, the MDP tool can accurately measure an upper bound for the security threshold. The results of our implementation show that the Bitcoin security threshold under our simplified model is lower than the security threshold obtained under the non-predictable WeRLman model.
@@
In the next step, we analyze the profitability of mining strategies before difficulty adjustment. An interesting observation under the volatile reward model is that a deviant strategy, such as a selfish mining attack, can become profitable immediately after launch without needing to wait for difficulty adjustment. As discussed in the literature (Grunspan and Pérez-Marco, 2018; Negy et al., 2020; Grunspan and Pérez-Marco, 2023; Sarenche et al., 2024a), under the fixed reward model assumption, a selfish miner incurs losses during the initial epoch of attack when the mining difficulty has not yet adjusted, the duration of which is at least two weeks. This initial loss period is one of the main reasons such attacks have not been observed in the Bitcoin network. However, under the volatile reward model, this loss period is shortened or may disappear entirely, making selfish mining a more significant threat in the transaction-fee era. Using our MDP-based tool within the simplified model, we analyze the Bitcoin security threshold prior to difficulty adjustment, which has not been explored in the WeRLman paper. Figure 1 illustrates the Bitcoin security threshold in the upcoming year as Bitcoin undergoes more halvings, based on different adversarial communication capabilities [^1] Communication capability is defined as the ratio of honest miners who receive the adversarial block first in the case of a block race.. Miners with a mining share above this threshold are not only incentivized to deviate from the honest strategy, but also do not face any initial loss. For a typical communication capability of $50\%$, when the protocol reward approaches 0, the Bitcoin security threshold reduces to $16\%$, which is within the range of current mining pool shares.
Figure 1. Bitcoin security threshold before DAM
@@ 2.1. System Model and Definitions
The honest strategy in Bitcoin protocol is defined as follows:
@@
In our model, time is assumed to be divided into rounds, denoted by $r$. In each round, a mining node can calculate multiple mining (hash) queries, the number of which is proportional to its mining share. A time step, denoted by $t$, represents the set of consecutive rounds after which a block is mined in the system. At each time step $t$, a block is generated by either the adversary or the honest miners. However, depending on the adversary’s strategy, the block generated at time step $t$ might not be immediately added to the canonical chain. Instead, it could be either added to the canonical chain or eventually be identified as an orphan block (excluded from the canonical chain) at a future time step.
@@
(1) $\begin{split}&\texttt{BR}^{\mathcal{A}}_{T}(\pi)=\frac{\sum_{t=0}^{T}{N_{\mathcal{A}}(t;\pi)}}{\sum_{t=0}^{T}{N_{\mathcal{A}}(t;\pi)+N_{\mathcal{H}}(t;\pi)}},\;\;\;\;\text{and}\;\;\;\;\\ &\texttt{BR}^{\mathcal{A}}(\pi)=\lim_{T\to\infty}{\texttt{BR}^{\mathcal{A}}_{T}(\pi)},\end{split}$
@@ 2.2. Overview of Selfish Mining Profitability
In Bitcoin, a difficulty epoch is defined as the period during which 2016 blocks are appended to the canonical chain. At the end of each epoch, the Difficulty Adjustment Mechanism (DAM) recalibrates the difficulty target for the subsequent epoch based on the hash power estimated from the previous epoch. Adversarial behaviors that destroy a portion of network’s mining power can cause DAM to underestimate the active mining power. This underestimation lowers mining difficulty, potentially making such behaviors profitable. In the following, we briefly review the well-established fact that, under the fixed reward model, adversarial behaviors like selfish mining are unprofitable before difficulty adjustment (Grunspan and Pérez-Marco, 2018; Sarenche et al., 2024b; Negy et al., 2020). However, they can become profitable over the long term after applying DAM.
2.3. Non-profitability of Pre-DAM Selfish Mining Under the Fixed Block Reward Model
@@
2.4. Profitability of Long-Range Selfish Mining Under the Fixed Block Reward Model

### v3

@@ Abstract.
In this paper, we present a reinforcement learning-based tool for analyzing mining strategies under a more realistic volatile reward model. The tool leverages the Asynchronous Advantage Actor-Critic (A3C) algorithm to derive near-optimal strategies while interacting with an environment that simulates the behavior of the Bitcoin mempool during any specified period, enabling analysis based on actual historical patterns. It supports the evaluation of adversarial mining strategies, such as selfish mining and undercutting, both before and after the difficulty adjustment, offering insights into the effects of mining attacks in both the short and long term.
@@ 1. Introduction
We proceed by introducing a simplified model to analyze mining strategies, in which the volatility of block rewards arises from the varying generation times of blocks. We argue that block reward volatility is not solely attributed to the infrequent arrival of whale transactions. The inherent competition among transactions to be included in the chain quickly causes blocks with longer generation times to include higher transaction fees, resulting in block reward volatility even in the absence of whale transactions. To this end, we build our simplified model upon the observed positive correlation between block generation time and its fee. Compared to the WeRLman environment, our simplified environment includes a greater number of blocks with varying fee levels, allowing us to analyze more fine-grained strategies. We employ a Markov Decision Process (MDP)-based tool to analyze the profitability of mining strategies under our simplified volatile model. Our MDP tool estimates a more accurate security threshold for Bitcoin in the transaction-fee era compared to the non-predictable WeRLman model. However, this threshold serves as an upper bound on the actual value, primarily due to the limited number of states the MDP can handle and the restriction of decision-making to block mining events.
@@
In the next step, we analyze the profitability of mining strategies before difficulty adjustment. An interesting observation under the volatile reward model is that a deviant strategy, such as a selfish mining attack, can become profitable immediately after launch without needing to wait for difficulty adjustment. As discussed in the literature (Grunspan and Pérez-Marco, 2018; Negy et al., 2020; Grunspan and Pérez-Marco, 2023; Sarenche et al., 2024a), under the fixed reward model assumption, a selfish miner incurs losses during the initial epoch of attack when the mining difficulty has not yet adjusted, the duration of which is at least two weeks. However, under the volatile reward model, this loss period is shortened or may disappear entirely, making selfish mining a more significant threat in the transaction-fee era. Using our MDP-based tool within the simplified model, we analyze the Bitcoin security threshold prior to difficulty adjustment, which has not been explored in the WeRLman paper. Figure 1 illustrates the Bitcoin security threshold in the upcoming year as Bitcoin undergoes more halvings, based on different adversarial communication capabilities [^1] Communication capability is defined as the ratio of honest miners who receive the adversarial block first in the case of a block race.. Miners with a mining share above this threshold are not only incentivized to deviate from the honest strategy, but also do not face any initial loss. For a typical communication capability of $50\%$, when the protocol reward approaches 0, the Bitcoin security threshold reduces to $16\%$, which is within the range of current mining pool shares.
Figure 1. Bitcoin security threshold before DAM.
@@ 2.1. System Model and Definitions
The honest strategy in the Bitcoin protocol is defined as follows:
@@
In our model, time is assumed to be divided into rounds, denoted by $r$. In each round, a mining node can calculate multiple mining (hash) queries, the number of which is proportional to its mining share. A time step, denoted by $t$, represents the set of consecutive rounds after which a block is mined in the system. At each time step $t$, a block is generated by either the adversary or the honest miners. However, depending on the adversary’s strategy, the block generated at time step $t$ might not be immediately added to the canonical chain. Instead, it could be either added to the canonical chain or eventually identified as an orphan block (excluded from the canonical chain) at a future time step.
@@
(1) $\begin{split}&\texttt{BR}^{\mathcal{A}}_{T}(\pi)=\frac{\sum_{t=0}^{T}{N_{\mathcal{A}}(t;\pi)}}{\sum_{t=0}^{T}{N_{\mathcal{A}}(t;\pi)+N_{\mathcal{H}}(t;\pi)}},\;\;\;\;\text{and}\;\;\;\;\\ &\texttt{BR}^{\mathcal{A}}(\pi)=\lim_{T\to\infty}{\texttt{BR}^{\mathcal{A}}_{T}(\pi)}\enspace,\end{split}$
@@ 2.2. Overview of Selfish Mining Profitability
In Bitcoin, a difficulty epoch is defined as the period during which 2016 blocks are appended to the canonical chain. At the end of each epoch, the Difficulty Adjustment Mechanism (DAM) recalibrates the difficulty target for the subsequent epoch based on the hash power estimated from the previous epoch. Adversarial behaviors that destroy a portion of the network’s mining power can cause DAM to underestimate the active mining power. This underestimation lowers mining difficulty, potentially making such behaviors profitable. In the following, we briefly review the well-established fact that, under the fixed reward model, adversarial behaviors like selfish mining are unprofitable before difficulty adjustment (Grunspan and Pérez-Marco, 2018; Sarenche et al., 2024b; Negy et al., 2020). However, they can become profitable over the long term after applying DAM.
2.2.1. Non-profitability of Pre-DAM Selfish Mining Under the Fixed Block Reward Model
@@
2.2.2. Profitability of Long-Range Selfish Mining Under the Fixed Block Reward Model